Lemma
EasyCounting & chance
OpenStax, Statistics §3.3 Two Basic Rules of Probability · Example 17 · Probability · Counting & chance

Two Basic Rules of Probability · Example 17

Refer to the information in. P = tests positive.

(a) Given that a woman develops the disease, what is the probability that she tests positive? Find P(P|B) = 1 − P(N|B).
(b) What is the probability that a woman develops the disease and tests positive? Find P(B AND P) = P(P|B)P(B).
(c) What is the probability that a woman does not develop the disease? Find P(B′) = 1 − P(B).
(d) What is the probability that a woman tests positive for the disease? Find P(P) = 1 − P(N).

~5 minWritten answer · compare with the solution

OpenStax, Statistics · §3.3 Two Basic Rules of Probability · Example 17

Stuck? A gentle nudge

No hints for this one.

Something off?
Workspace
Your working

Check your full solution

Upload photos of your handwritten work or a PDF, and Lemma will read it and tell you whether it's right — and where it first goes wrong if not.